OpenStudy (anonymous):

I really need help, if you know how to do this and can it explain how you got an answer that would be great cause I have no clue how to do this problem NO2 is known to decompose via a second-order mechanism. If the half-life of this decomposition is 250 seconds when the initial concentration of NO2 is 0.75 M, what is the rate constant for this decomposition? A) 0.0053 s-1 B) 190 s-1 C) 0.0027 s-1 D) 360 s-1

OpenStudy (aaronq):

do you know the 2nd order rate law (formula)?

OpenStudy (anonymous):

yes there is two 1/ [A]t - 1/[A]o =kt and then t1/2 =1/k[A]o i have tried using them but what throws me off is that they give and initial rate but no final and that's when i get lost

OpenStudy (aaronq):

use the second formula you posted. They didn't give you any rates.

OpenStudy (aaronq):

\(t_{1/2}=\dfrac{1}{k[A]_o}\)